Suppose you need to construct a right triangle in which the shorter side is eight feet less than the longest side and the third side is seven feet more than the shortest side.
Answer by checkley77(12844)

a^2+b^2=c^2
a=c-8
b=a+7 or b=(c-8)+7 or b=c-1
(c-8)^2+(c-1)^2=c^2
c^2-16c+64+c^2-2c+1=c^2
2c^2-c^2-18c+65=0
c^2-18c+65=0
(c-13)(c-5)=0
c-13=0
c=13 ans.
a=13-8
a=5 ans.
b=5+7
b=12 ans.
Proof:
5^2+12*2=13^2
25+141=169
169=169